# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ Abstract
Mobile health leverages personalized, contextually-tailored interventions optimized through bandit and reinforcement learning algorithms. Despite its promise, challenges like participant heterogeneity, nonstationarity, and nonlinearity in rewards hinder algorithm performance. We propose a robust contextual bandit algorithm, termed “DML-TS-NNR”, that simultaneously addresses these challenges via (1) modeling the differential reward with user- and time-specific incidental parameters, (2) network cohesion penalties, and (3) debiased machine learning for flexible estimation of baseline rewards. We establish a high-probability regret bound that depends solely on the dimension of the differential reward model. This feature enables us to achieve robust regret bounds even when the baseline reward is highly complex. We demonstrate the superior performance of the DML-TS-NNR algorithm in a simulation and two off-policy evaluation studies.
@@ 1 Introduction
Mobile health (mHealth) uses smart devices to deliver digital notification interventions to users. These notifications nudge users toward healthier attitudes and behaviors. Because mHealth applications can monitor and react to users and their environment, they offer the promise of personalized, contextually-tailored interventions. In practice, achieving this promise requires bandit or reinforcement learning (RL) algorithms that can accurately learn mHealth intervention effects, including how effects vary by user, over time, and based on context. In this regard, contextual bandit algorithms are appealing due to strong empirical performance in other settings, their ability to customize intervention decisions based on changing contexts, and their simplicity and extensibility relative to full RL algorithms.
One unique aspect of mHealth (Greenewald et al., 2017) is the presence of a “do-nothing” action in which the mHealth app does not intervene. As mHealth rewards typically involve human decision making (e.g., whether the user chooses to exercise), the distribution of the the corresponding baseline reward—the reward under the “do-nothing” action—typically evolves in a complex, nonstationary fashion as time progresses and contexts change. In contrast, the treatment effects—the difference in the expected value of the reward relative to the “do-nothing” action—tend to be much more stable and can be adequately modeled using classical stationary models; e.g., linear models (Robinson, 1988).
Greenewald et al. (2017) introduced an action-centered contextual bandit algorithm with sub-linear regret in this setting by replacing the observed reward, $R_{t}$, with a “pseudo-reward,” $\tilde{R}_{t}$, where $t$ indexes the time points. Denoting the binary action as $A_{t}\in\{0,1\}$ with $A_{t}=0$ corresponding to the “do-nothing” action, the pseudo-reward for $A_{t}$ can be written as
$\tilde{R}_{t}:=(A_{t}-\pi_{t})R_{t}=\pi_{t}(1-\pi_{t})R_{t}\left(\frac{A_{t}}{\pi_{t}}-\frac{1-A_{t}}{1-\pi_{t}}\right),$
where $\pi_{t}$ is the (known) probability of treatment assignment from the Thompson-sampling algorithm. We took an additional step to derive an equivalent expression (second equality), which reveals that $\tilde{R}_{t}$ is proportional to an inverse-probability-weighted (IPW) estimator. Crucially, $\tilde{R}_{t}$ provides an unbiased estimate of the treatment effect regardless of the non-stationarity of the baseline outcome.
Although Greenewald et al. (2017)’s analysis shows sub-linear regret in nonstationary settings, our simulations show that it is outperformed by other methods over finite time horizons. The reasons include its failure to address treatment effect heterogeneity, pool information across users, and model the baseline outcome, which leads to highly variable pseudo-rewards and slow learning. We propose to generalize their method to overcome these limitations by
@@
    imposing hierarchical structure in the treatment effect model with shared structural parameters and “incidental parameters” for users and time,
@@
    efficiently pooling across users and time via nearest-neighbor regularization (NNR),
@@
    de-noising rewards with flexible supervised learning models via the de-biased machine learning (DML) framework of Chernozhukov et al. (2018).
We name the resulting method DML-TS-NNR to highlight its usage of DML, NNR, and Thompson Sampling. The primary contributions are (a) the DML-TS-NNR method, (b) a high-probability regret bound that relies solely on the dimension of the differential reward model, which is typically much smaller than that of the complex baseline reward model, and (c) empirical comparisons demonstrating the superior performance of DML-TS-NNR in simulation and two real-world mHealth studies.
@@ 2 Related Work
Related work in statistics considers treatment effect estimation with longitudinal data. Cho et al. (2017) present a semiparametric random effects method for estimating subject-specific static treatment effects. Qian et al. (2020) discusses the statistical challenges of applying linear mixed-effects models to mHealth data and shows that the resulting estimates may lack a causal interpretation.
In the bandit literature, several works address a growing pool of users. Ma et al. (2011) introduced NNR in recommender systems using homophily principles—similar nodes are more likely to be connected than dissimilar ones (McPherson et al., 2001). Cesa-Bianchi et al. (2013) adapted NNR for bandit settings, improving regret. Subsequent improvements focused on scalability and algorithm modifications for stronger regret bounds (Vaswani et al., 2017; Yang et al., 2020).
@@
Aouali et al. (2023) and Lee et al. (2024) also propose bandit algorithms with mixed effects; methodologically, however, their use of mixed effects is dissimilar from that of the works above and our own. Aouali et al. (2023) employs random effects to relate treatments with categories (e.g., movies within a genre), and Lee et al. (2024) considers the case in which a single, discrete action leads to a multivariate reward that is affected by a user-specific random effect. In contrast, the other approaches listed above (and our own) do not assume any relationships among the actions, and they associate each discrete action with a single scalar-valued reward in an iterative fashion.
Other work develops semiparametric methods that lead to sublinear regret without requiring correct specification of the conditional mean reward. Krishnamurthy et al. (2018) and Kim & Paik (2019) propose improvements to the approach of Greenewald et al. (2017) that lead to stronger regret bounds. Kim et al. (2021, 2023) develop a doubly-robust approach for both linear and generalized linear contextual bandits. However, in contrast to Tomkins et al. (2021), Hu et al. (2021), and our approach, these semiparametric methods are not immediately applicable to longitudinal settings.
@@ 3.1 Problem Statement
Recognizing the connection between mHealth policy learning and contextual bandit methods, we now consider a contextual multi-armed bandit environment with one control arm ($a=0$) and $q$ non-baseline arms corresponding to different actions or treatments. Individuals are indexed by $i=1,2,\ldots$, and decision points are indexed by $t=1,2,\ldots$. For each individual $i$ at time $t$, the algorithm observes a context vector $S_{i,t}\in\mathcal{S}$, chooses an action $A_{i,t}\in[q]:=\{0,\ldots,q\}$, and receives a reward $R_{i,t}\in\mathbb{R}$. The differential reward describes the difference in expected rewards between choosing a non-baseline arm and the control arm, defined as $\Delta_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]-\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=0\right]$.
We denote the observation history up to time $t$ for individual $i$ by $\mathcal{H}_{i,t}:=(S_{i1},A_{i1},R_{i1},\dots,R_{it})$. Actions are selected according to stochastic policies, $\pi_{i,t}:\mathcal{H}_{i,t}\times\mathcal{S}\to\mathcal{P}([q])$. We use $\pi_{i,t}(a|s)$ to denote the probability of action $a\in[q]$ given current context $s\in\mathcal{S}$ for a fixed (implicit) history. As in Greenewald et al. (2017), we assume that the probability of control action is bounded:
@@
There exists $10<\pi_{\min},\pi_{\max}<$ 1 s.t $\pi_{\min}<\pi_{i,t}(0|a),1-\pi_{i,t}(0|a)<\pi_{\max}\forall i,t,a$.
As Greenewald et al. (2017) noted, this assignment probability controls the number of messages that participants receive. It ensures that participants receive sufficient messages that they do not disengage, but not so many that they are overwhelmed or fatigued. We now define
$A^{*}=\operatorname{argmax}_{a\in[q]}\mathbb{E}\left[R_{i,t}|S_{i,t},A_{i,t}=a\right],\quad\bar{A}^{*}=\operatorname{argmax}_{a\in[q]\backslash\{0\}}\Delta_{i,t}(S_{i,t},a),$
the optimal treatment arm (including control) and the alternative arm, respectively. Respecting the restrictions in Assumption 1, we can now express the optimal policy, $\pi_{i,t}^{*}$, in two cases:
@@
    $A^{*}>0$: $\pi_{i,t}^{*}(A^{*}|S_{i,t})=1-\pi_{\min}$ and $\pi_{i,t}^{*}(0|S_{i,t})=\pi_{\min}$.
These cases represent achieving the appropriate boundary condition given the optimal action, i.e., maximum chance of “do-nothing” action if $A^{\star}=0$ and minimum chance if $A^{\star}\not=0$. Otherwise, $\pi_{i,t}^{*}(a|S_{i,t})=0$. In Section 5.2, we define regret relative to this optimal policy, $\pi_{i,t}^{*}$.
Lastly, we consider a study design that progresses in stages, where individuals enter the study sequentially (Friedman et al., 2010), as illustrated in Figure 3 in Appendix A. Staged recruitment serves two purposes in our analysis: it increases the fidelity of the theoretical setting by mimicking real-world recruitment strategies, and it allows us to estimate changes in the differential reward over time. Concretely, we first observe individual $i=1$ at time $t=1$. Then we observe individuals $i=(1,2)$ at times $t=(2,1)$. Subsequent stages $k$ involve individuals $i\leq k$, each having had $k-i+1$ observations, respectively. Define $\mathcal{O}_{k}=\{(i,t):i\leq k\>\&\>t\leq k+1-i\}$ as the double index set of all observations in stage $k$. A notation summary is presented in Appendix G.
@@ 3.2 Doubly Robust Differential Reward
Following Greenewald et al. (2017), we assume that the differential reward is linear:
$\Delta_{i,t}(s,a)=x(s,a)^{\top}\theta_{i,t},$ (1)
where $x(s,a)\in\mathbb{R}^{p\times 1}$ is a feature vector depending on the context and action. We allow the baseline reward, $g_{t}(s)$, to be a complex, nonlinear, and nonstationary function of context and time. Combined, these two assumptions imply a partially linear model for the conditional mean reward:
$\begin{split}\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]&=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s),\end{split}$ (2)
@@
$R_{i,t}=x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+g_{t}(s)+\epsilon_{i,t},$
where we make the following standard (Abeille & Lazaric, 2017) assumption regarding $\epsilon_{i,t}$:
@@
$\epsilon_{i,t}$ is conditionally mean zero (i.e., $\mathbb{E}[\epsilon_{i,t}|\mathcal{H}_{i,t}]=0$) and sub-Gaussian with variance $\sigma^{2}$: $\mathbb{E}[\exp(\eta\epsilon_{i,t})|\mathcal{H}_{i,t}]\leq\exp(\eta^{2}\sigma^{2}/2)$ for $\eta>0$.
Under this model assumption, the IPW estimator of Greenewald et al. (2017) produces an unbiased estimate of $\Delta_{i,t}(s,a)$; however, it suffers from high variance, which leads to suboptimal performance in practice. Our proposed algorithm DML-TS-NNR, however, relies on an improved estimator, $\tilde{R}_{i,t}^{f}(s,\bar{a})$, that is also unbiased but achieves lower variance by denoising the rewards according to a working model, $f_{i,t}(s,a)$, for the conditional mean $r_{i,t}(s,a):=\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$:
@@
where $\bar{a}$ is a pre-selected non-baseline arm; i.e., we condition on $A_{i,t}\in\{0,\bar{a}\}$. The first component is a model-based prediction of the differential reward, and the second component is a debiasing term that uses IPW to correct for potential errors in specifying $f$. This improved estimator is a form of pseudo-outcome, a common technique in the causal inference literature (Bang & Robins, 2005; Kennedy, 2020; Nie & Wager, 2021). We refer to it as a Doubly Robust Differential Reward because it produces an unbiased estimate of $\Delta_{i,t}(s,a)$ as long as either $\pi_{i,t}$ or $f_{i,t}$ is correctly specified (see proof in Appendix F.1); though, in practice, the benefit stems from (a) the robustness to misspecification of $f$ (because $\pi_{i,t}$ is known) and (b) the variance reduction introduced by de-noising the outcomes. Lemma 5 and Remark 2 in Appendix F.2 provide a proof and discussion of the variance reduction. In the following, we abbreviate $\tilde{R}_{i,t}^{f}(s,a)$ as $\tilde{R}_{i,t}^{f}$, with the state and action implied.

### v4

# RoME: A Robust Mixed-Effects Bandit Algorithm for Optimizing Mobile Health Interventions

@@ Abstract
Mobile health leverages personalized and contextually tailored interventions optimized through bandit and reinforcement learning algorithms. In practice, however, challenges such as participant heterogeneity, nonstationarity, and nonlinear relationships hinder algorithm performance. We propose RoME, a Robust Mixed-Effects contextual bandit algorithm that simultaneously addresses these challenges via (1) modeling the differential reward with user- and time-specific random effects, (2) network cohesion penalties, and (3) debiased machine learning for flexible estimation of baseline rewards. We establish a high-probability regret bound that depends solely on the dimension of the differential-reward model, enabling us to achieve robust regret bounds even when the baseline reward is highly complex. We demonstrate the superior performance of the RoME algorithm in a simulation and two off-policy evaluation studies.
@@ 1 Introduction
Mobile health (mHealth) uses smart devices to deliver digital notification interventions to users. These notifications nudge users toward healthier attitudes and behaviors. Because mHealth applications can monitor and react to users and their environment, they offer the promise of personalized, contextually tailored interventions. In practice, achieving this promise requires bandit or reinforcement learning (RL) algorithms that can accurately learn mHealth intervention effects, including how they vary by user, over time, and based on context. In this regard, contextual bandit algorithms are appealing due to strong empirical performance in other settings, their ability to customize intervention decisions based on changing contexts, and their simplicity and extensibility relative to full RL algorithms.
One unique aspect of mHealth (Greenewald et al., 2017) is the presence of a “do-nothing” action in which the mHealth app does not intervene. As mHealth rewards typically involve human decision making (e.g., whether the user chooses to exercise), the distribution of the corresponding baseline reward—the reward under the do-nothing action—typically evolves in a complex, nonstationary fashion as time progresses and contexts change. In contrast, the treatment effects—the difference in the expected value of the reward relative to the do-nothing action—tend to be much more stable and can be adequately modeled using classical stationary models, such as linear models (Robinson, 1988).
Greenewald et al. (2017) introduced an action-centered contextual bandit algorithm with sublinear regret in this setting by replacing the observed reward, $R_{t}$, with a “pseudo-reward,” $\tilde{R}_{t}$, where $t$ indexes the time points. Denoting the binary action as $A_{t}\in\{0,1\}$ with $A_{t}=0$ corresponding to the do-nothing action, the pseudo-reward for $A_{t}$ can be written as
$\tilde{R}_{t}\coloneq(A_{t}-\pi_{t})R_{t}=\pi_{t}(1-\pi_{t})R_{t}\left(\frac{A_{t}}{\pi_{t}}-\frac{1-A_{t}}{1-\pi_{t}}\right),$
where $\pi_{t}$ is the (known) probability of treatment assignment from the Thompson-sampling algorithm. We took an additional step to derive an equivalent expression (second equality), which reveals that $\tilde{R}_{t}$ is proportional to an inverse-probability-weighted (IPW) estimator. Crucially, $\tilde{R}_{t}$ provides an unbiased estimate of the treatment effect regardless of the nonstationarity of the baseline outcome.
Although Greenewald et al. (2017)’s analysis shows sublinear regret in nonstationary settings, our simulations show that other methods can achieve lower regret over finite time horizons. The reasons include its failure to address treatment effect heterogeneity, pool information across users, and model the baseline outcome, which leads to highly variable pseudo-rewards and slow learning. We propose to generalize their method to overcome these limitations by
@@
    imposing hierarchical structure in the treatment effect model with shared fixed effects and random effects for users and time (i.e., a mixed-effects model),
@@
    efficiently pooling across users and time via nearest-neighbor regularization (NNR), and
@@
    denoising rewards with flexible supervised learning models via the debiased machine learning (DML) framework of Chernozhukov et al. (2018).
We name the resulting method RoME to highlight that it is a Robust Mixed-Effects algorithm. The primary contributions are (a) the RoME method, (b) a high-probability regret bound that relies solely on the dimension of the differential-reward model, which is typically much smaller than that of the complex baseline reward model, and (c) empirical comparisons demonstrating the superior performance of RoME in simulation and two real-world mHealth studies.
@@ 2 Related Work
Related work in statistics considers the estimation of treatment effects with longitudinal data. Cho et al. (2017) present a semiparametric random-effects method for estimating subject-specific static treatment effects. Qian et al. (2020) discusses the statistical challenges of applying linear mixed-effects models to mHealth data and shows that the resulting estimates may lack a causal interpretation.
In the bandit literature, several works address the challenge of heterogeneous users. Ma et al. (2011) introduced NNR in recommender systems using homophily principles—similar nodes are more likely to be connected than dissimilar ones (McPherson et al., 2001). Cesa-Bianchi et al. (2013) adapted NNR for bandit settings, improving regret. Subsequent improvements focused on scalability and algorithm modifications for stronger regret bounds (Vaswani et al., 2017; Yang et al., 2020).
@@
Aouali et al. (2023) and Lee et al. (2024) also propose bandit algorithms with mixed effects; methodologically, however, their use of mixed effects differs from the above approaches and our own. Aouali et al. (2023) employs random effects to relate treatments within categories (e.g., movies within a genre), and Lee et al. (2024) considers the case in which a single, discrete action leads to a multivariate reward that is affected by a user-specific random effect. In contrast, the other approaches listed above (and our own) do not assume any relationships among the actions, and they associate each discrete action with a single scalar-valued reward in an iterative fashion.
Other work develops semiparametric methods that lead to sublinear regret without requiring correct specification of the conditional mean reward. Krishnamurthy et al. (2018) and Kim & Paik (2019) propose improvements to the approach of Greenewald et al. (2017) that lead to stronger regret bounds. Kim et al. (2021, 2023) develop a doubly robust approach for both linear and generalized linear contextual bandits. However, in contrast to Tomkins et al. (2021), Hu et al. (2021), and our approach, these semiparametric methods are not immediately applicable to longitudinal settings.
@@ 3.1 Problem Statement
Recognizing the connection between mHealth policy learning and contextual bandit methods, we now consider a contextual multi-armed bandit environment with one control arm (the do-nothing action) denoted by $a=0$ and $q$ non-baseline arms corresponding to different actions or treatments. Users are indexed by $i=1,2,\ldots$ and decision points are indexed by $t=1,2,\ldots$. For each user $i$ at time $t$, the algorithm observes a context vector $S_{i,t}\in\mathcal{S}$, chooses an action $A_{i,t}\in[q]\coloneq\{0,\ldots,q\}$, and receives a reward $R_{i,t}\in\mathbb{R}$. The differential reward describes the difference in expected rewards between choosing a non-baseline arm and the control arm, defined as $\Delta_{i,t}(s,a)\coloneq\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]-\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=0\right]$.
We denote the observation history up to time $t$ for user $i$ as $\mathcal{H}_{i,t}\coloneq(S_{i1},A_{i1},R_{i1},\dots,S_{i,t})$. Actions are selected according to stochastic policies, $\pi_{i,t}:\mathcal{H}_{i,t}\times\mathcal{S}\to\mathcal{P}([q])$. We use $\pi_{i,t}(a|s)$ to denote the probability of action $a\in[q]$ given context $s\in\mathcal{S}$ for a fixed (implicit) history. As in Greenewald et al. (2017), we assume that the probability of the control action is bounded:
@@
There exists $10<\pi_{\min},\pi_{\max}<$ 1 such that $\pi_{\min}<\pi_{i,t}(0|a)<\pi_{\max}$ for all $i,t,a$.
As noted in Greenewald et al. (2017), this assignment probability controls the number of messages that participants receive. It ensures that participants receive sufficient messages that they do not disengage, but not so many that they are overwhelmed or fatigued. We now define
$A_{i,t}^{\star}=\underset{a\in[q]}{\operatorname{argmax}}\ \mathbb{E}\left[R_{i,t}|S_{i,t},A_{i,t}=a\right],\quad\bar{A}_{i,t}^{\star}=\underset{a\in[q]\backslash\{0\}}{\operatorname{argmax}}\ \mathbb{E}\left[R_{i,t}|S_{i,t},A_{i,t}=a\right],$
the optimal arm (including control) and the optimal non-baseline arm, respectively. Respecting the restrictions in Assumption 1, we can now express the optimal policy, $\pi_{i,t}^{\star}$, in two cases:
@@
    $A_{i,t}^{\star}>0$: $\pi_{i,t}^{\star}(0|S_{i,t})=\pi_{\min}$ and $\pi_{i,t}^{\star}(\bar{A}_{i,t}^{\star}|S_{i,t})=\pi_{i,t}^{\star}(A_{i,t}^{\star}|S_{i,t})=1-\pi_{\min}$.
These cases represent the appropriate boundary condition given the optimal action. We maximize the probability of the control action if $A_{i,t}^{\star}=0$ and minimize it if $A_{i,t}^{\star}\not=0$, selecting the next-best action with the highest possible probability; we set the assignment probabilities for the remaining arms to zero. In Section 5.2, we define regret relative to this optimal policy, $\pi_{i,t}^{\star}$.
Figure 1: Illustration of the staged recruitment scheme. At each recruitment stage (each time point), a new participant is recruited and observed. At the same time, all participants who were recruited prior to the current stage are also observed again. Observations are not collected from participants who have yet to be recruited. For simplicity, we assume one participant is recruited at each stage.
Lastly, we consider a study design that progresses in stages, where users enter the study sequentially (Friedman et al., 2010), as illustrated in Figure 1. Staged recruitment serves two purposes in our analysis: it increases the fidelity of the theoretical setting by mimicking real-world recruitment strategies, and it allows us to estimate changes in the differential reward over time. Concretely, we first observe user $i=1$ at time $t=1$. Then we observe users $i=(1,2)$ at times $t=(2,1)$. Subsequent stages $k$ involve users $i\leq k$, each having had $k-i+1$ observations, respectively. Define $\mathcal{O}_{k}=\{(i,t):i\leq k,\,t\leq k+1-i\}$ as the set of all observations in stage $k$. Appendix F provides a notation summary.
@@ 3.2 Doubly Robust Differential Reward
Following Greenewald et al. (2017), we assume that the differential reward is linear: $\Delta_{i,t}(s,a)=x(s,a)^{\top}\theta_{i,t}$, where $x(s,a)\in\mathbb{R}^{d}$ is a feature vector depending on the context and action. We allow the baseline reward, $g_{t}(s)$, to be a nonlinear and nonstationary function of context and time. Combined, these two assumptions imply a partially linear model for the conditional mean reward:
$\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]=g_{t}(s)+x(s,a)^{\top}\theta_{i,t}\delta_{a>0},$
@@
$R_{i,t}=g_{t}(s)+x(s,a)^{\top}\theta_{i,t}\delta_{a>0}+\epsilon_{i,t},$
where we make the following standard assumption (Abeille & Lazaric, 2017) regarding $\epsilon_{i,t}$:
@@
The error $\epsilon_{i,t}$ is conditionally mean zero (i.e., $\mathbb{E}[\epsilon_{i,t}|\mathcal{H}_{i,t}]=0$) and sub-Gaussian.
Assumption 2 ensures that large errors are sufficiently rare and that $\tilde{R}_{t}$ is proportional to an unbiased estimate of $\Delta_{i,t}(s,a)$. However, $\tilde{R}_{t}$ suffers from high variance, which leads to suboptimal performance in practice. Our proposed algorithm, RoME, relies on an improved estimator, $\tilde{R}_{i,t}^{f}(s,\bar{a})$, that is also unbiased but achieves lower variance by denoising the rewards according to a working model, $f_{i,t}(s,a)$, for the conditional mean $r_{i,t}(s,a)\coloneq\mathbb{E}\left[R_{i,t}|S_{i,t}=s,A_{i,t}=a\right]$:
@@
where $\bar{a}$ is a preselected non-baseline arm; i.e., we condition on $A_{i,t}\in\{0,\bar{a}\}$. The first component is a model-based prediction of the differential reward, and the second component is a debiasing term that uses IPW to correct for potential errors in the specification of $f$. This improved estimator is a form of pseudo-outcome, a common technique in the causal inference literature (Bang & Robins, 2005; Nie & Wager, 2021; Kennedy, 2023). We refer to it as a Doubly Robust Differential Reward because it produces an unbiased estimate of $\Delta_{i,t}(s,a)$ as long as either $\pi_{i,t}$ or $f_{i,t}$ is correctly specified (see proof in Appendix E.1); though, in practice, the benefit stems from (a) the robustness to misspecification of $f$ (because $\pi_{i,t}$ is known) and (b) the variance reduction introduced by denoising the outcomes (see Remark 1 in Appendix E.2). In the following, we abbreviate $\tilde{R}_{i,t}^{f}(s,a)$ as $\tilde{R}_{i,t}^{f}$, with the context and action implied.